Shutdown activities

Infrastructure

TASK Build. DATE DOCUMENTS PERSONNEL COMMENTS- IMPACT STATUS
Ventillation monitoring Safe Area PZ85 05/02/2010   G. Morard No impact Ongoing
HVAC units for D3 A and D3 B UX85A and B Weeks 2 and 3 AOC 4251907 VIC 14/12 EN/CV (+Initiative + Cerclim) - EN/EL - ALTEAD Shielding Wall Mobile blug open from 04/01 Greene nacelle needed A side Ongoing
SG8 / RICH1 cooling / Buffer volume SG8 Week 3 or 4 (TBC)   EN/CV (+Initiative ) No mixed water in SG for 1 day Postoned to 01/03/2010
Detector Cooling - Maintenance (UPS) UX85A 04/01   EN/CV F. Corbaz All 4 Detector cooling plants off Done
Detector Cooling - Consolidation (Hydraulic ) UX85A TBD   EN/CV F. Corbaz All 4 Detector coling plants off POSTPONED
Detector Cooling - Maintenance (Hydraulic - Leaking valve ) UX85A TBD   EN/CV F. Corbaz RICH2 off Done
Rich 1 and 2 gas system maintenance (membrane + recovery + purif) SG8 TBD   PH/DT S.Pavis 1h unavailibility during mambranne replacment Done
RAMSES probes in extraction ducts SX8 01/02/2010   EN/ CV G. Morard No air extraction in UX and SX (Limited or No work in UX!) Done
CV primary cooling maintenance and consolidation: SF8 by-pass + Trane 801,802 and 803 P8 Week 7   EN/ CV G. Morard No redudency in primary cooling for ~1 w Ongoing
Change of communication card on ventillation PLC SUX8 Week 6   G. Morard Stop of ventillation for 10'. To be scheduled before 08:00 or after 18:00

Safety Systems

TASK Build. DATE DOCUMENTS PERSONNEL COMMENTS STATUS
CO2 system Calibration and maint. UX85A, UX85B Week 5 EDMS L. Roy On site work Done
Water Mist valve protection UX85B Week 4   PH - LBO B. Chadaj Inhibit DSS and Close Safety valve on magnet Done
IS37 test of ODH and EVAC UX85 22/12/2009   GS/SEM D. Raffour No access to UX85 Done
PZ85 latency UX85 23/12/2009   EN/HE JP Lappe No access to UX85 Done
Flammable gas Cabling SG8 05/01   EN/EL L. Roy Gas system stops (1h am and 10' pm) Done
RAMSES installation in air exhaust ducts UX85 01/02   EN/CV G.Morard Ventillation degraded in UX85. No work underground on 01/02
SNIFFER IS37 maintenance and test UX85A 8-12/02   GS/ASE D. Raffourt System not available 2h for each line - detector

-- EricThomas - 04-Jan-2010

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2010-02-09 - EricThomas
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback